OpenStudy (anonymous):

QUICK HELP PLEASE !! MEDAL AND FAN Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y = 4x - 11 y = 4x + 11 y = 4x + 7 y = 4x - 7

OpenStudy (cydney_morgan):

Do you know how to do this Bria?

OpenStudy (anonymous):

no its a Pre-test

OpenStudy (cydney_morgan):

Oh, I see, I'll get one of my friends to come explain to you how to do this ok? I don't know but I'll hook you up.

OpenStudy (cydney_morgan):

@iambatman

OpenStudy (cydney_morgan):

@Squirrels

OpenStudy (anonymous):

Thanks !

OpenStudy (cydney_morgan):

@undeadknight26

OpenStudy (anonymous):

You want to set it up as y =mx+b form where m is your slope

OpenStudy (anonymous):

Do you know of it?

OpenStudy (anonymous):

y=4x , thats it how do i get B

OpenStudy (anonymous):

Well first thing first, we have to use point - slope form. \[\huge y -y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

I know this

OpenStudy (anonymous):

Ok so what's the problem then?

OpenStudy (anonymous):

You have (x1,y1), do some algebra, and boom you're done.

OpenStudy (anonymous):

Okay thanks for the help i got it from here

OpenStudy (anonymous):

Alright, good luck.